# Model-independent search for T violation with T2HK and DUNE

### v1

@@
August 12, 2024
@@ 2.1 Framework and assumptions
$\phi_{ij}\approx\frac{\Delta m^{2}_{ij,{\rm eff}}(E_{\nu})L}{2E_{\nu}}\,,$ (2.11)
@@ 3.1 Experimental setups
DUNE is an upcoming long-baseline accelerator neutrino experiment designed to explore the nature of neutrinos by sending them from Fermilab to a far-detector situated deep underground at the Sanford Lab in South Dakota at a distance of 1300 km. To estimate the experiment’s sensitivity, we utilize the configurations outlined in the Technical Design Report (TDR) [25, 34], along with an alternative configuration featuring enhanced energy resolution as suggested in Refs. [35, 36] (see also [37, 38]). The TDR configuration includes a 40 kt Liquid Argon Time Projection Chamber (LArTPC) as the far detector (current planning considers a staged approach with up to 4 detector modules of 17 kt each), paired with a 120 GeV proton beam delivering 1.2 MW of beam power, which translates to $1.1\times 10^{21}$ protons on target (P.O.T) per year. For further details on systematic errors and efficiencies, please refer to Ref. [34].
T2HK (Tokai to Hyper-Kamiokande) is an off-axis, accelerator-based future superbeam experiment with a 295 km baseline. To estimate the detector’s physics potential, we adhere to the experimental configurations outlined in [24]. This experiment will utilize the same 30 GeV proton beam from the J-PARC facility, previously used for T2K, to generate (anti)neutrino fluxes. The Water Cherenkov far detector is expected to have a fiducial volume of 187 kt, and the total exposure will be $1.3\,\rm{MW}\times 10\times 10^{7}$ seconds, equivalent to $2.7\times 10^{22}$ protons on target (P.O.T). In this simplified scenario, we consider an uncorrelated 5% (3.5%) signal normalization error, a 10% background normalization error, and a 5% energy calibration error for both $\nu$ and $\bar{\nu}$ appearance (disappearance) channels.
@@ 3.2 Analysis details
Here $k$ labels the energy bin and $N_{k}^{x}(c_{i},\xi_{x})$ is the number of events predicted in the model eq. 2.3 for experiment $x={\rm T2HK,DUNE}$, calculated including backgrounds and various systematics as described in section 3.1. The latter are parametrised by pull parameters, generically denoted by $\xi_{x}$ in eq. 3.1. $N_{k}^{x,\rm obs}$ is the corresponding “observed” number of events, which will depend on the true mechanism of neutrino conversion realised in Nature. In this study we will—as a specific example—always assume the standard three-flavour scenario and calculate $N_{k}^{x,\rm obs}$ accordingly, using the oscillation parameters shown in table 1. Then we study the sensitivity of the T-violation test as a function of the assumed true value of $\delta_{\rm CP}$, as $N_{k}^{x,\rm obs}(\delta_{\rm CP})$. Hence, $\chi^{2}_{k}(c_{i})\to\chi^{2}_{k}(c_{i},\delta_{\rm CP}^{\rm true})$. Obviously, there can only be sensitivity for T-violation for values $\delta_{\rm CP}\neq 0,\pi$. To calculate $N_{k}^{x,\rm obs}$ in the standard three-flavour case we use a line-averaged constant matter density of 2.84 g/$\rm{cm}^{3}$ [40, 41] for both, DUNE and T2HK, which is a good approximation for these baselines [23]. The same value is then adopted to calculate the oscillation frequencies in the T conserving model according to eq. 2.11.222In this approximation the model implemented by eq. 2.3 can reproduce event numbers for T conservation $\delta_{\rm CP}=0,\pi$ exactly.
@@ 4.1 Exposure and energy resolution
Let us now study the interplay of energy resolution and exposure. The estimates in section 3.3 suggest that the statistical uncertainty is dominated by DUNE under our default assumptions of 608 (840) kt MW yr for T2HK (DUNE), see eq. 3.5. Figure 4 shows the T violation sensitivity for $\delta_{\rm CP}^{\rm tr}=90^{\circ}$ as a function of the DUNE exposure in neutrino mode for different assumptions on the energy resolutions in both experiments. We see that an improved energy reconstruction can somewhat reduce the required exposure to reach a certain sensitivity. Note that our default assumption for DUNE according to eq. 4.1 corresponds already to an improved resolution according to [36]. The magenta curve in fig. 4 shows the sensitivity for a DUNE resolution reduced by a factor to compared to eq. 4.1, which corresponds to good accuracy to the value from the DUNE TDR [25, 34]. We see that the improved reconstruction is essential to obtain good sensitivity with reasonable exposures.
@@ 4.2 Zero-distance effect and prior on oscillation frequencies
Figure 5 shows a non-trivial interplay of the near-detector constrains (curves with different colors), the energy resolution (solid curves: default assumption, dashed curves: improved by a factor 0.5), and the oscillation frequency prior. A crucial role is played by the quantity $\delta_{0}$ defined in eq. 2.18, which multiplies the non-unitarity factor $\epsilon^{2}$ in the observable $X_{T}$, see eq. 2.17. The value of $\delta_{0}$ is determined by the oscillation frequencies and therefore depends on energy. Hence, effectively we have to average $\delta_{0}$ over the width of the considered energy bin and fold it with the resolution function. It turns out that the energy-averaged value of $\delta_{0}$ monotonically decreases from about $-0.09$ for our default energy resolution to $-0.21$ for a resolution improved by a factor 0.2. Hence, we expect that the impact of the zero-distance effect becomes more important for better energy resolution. This is the behaviour visible in fig. 5: while the solid curves are rather insensitive to the value of the near-detector constraint even up to $\sigma_{\epsilon}=5\%$, we see a rather strong dependence on $\sigma_{\epsilon}$ for the improved energy resolution (dashed curves) and the sensitivity degrades significantly for uncertainties $\sigma_{\epsilon}\gtrsim 1\%$. This behaviour is consistent with the estimate in eq. 3.5. Note that for $\sigma_{\epsilon}=5\%$ the better resolution leads to a worse sensitivity, which is a manifestation of the larger value of $|\delta_{0}|$ due to the non-trivial energy averaging, see also eq. 3.4.
@@ 5 Summary and discussion
[…]
@@ Appendix A Time reversal in the QFT formalism for neutrino oscillations
[…]
@@ Appendix B Constraints on the zero-distance effect
    where $N$ ($D$) is the numerator (denominator) of eq. B.3. In the limit of small $\sigma_{\epsilon,\rm ext}$ we obtain

### v2

@@
October 15, 2025
We consider the time reversal (T) transformation in neutrino oscillations in a model-independent way by comparing the observed transition probabilities at two different baselines at the same neutrino energy. We show that, under modest model assumptions, if the transition probability $P_{\nu_{\mu}\to\nu_{e}}$ around $E_{\nu}\simeq 0.86$ GeV measured at DUNE is smaller than the one at T2HK the T symmetry has to be violated. Experimental requirements needed to achieve good sensitivity to this test for T violation are to obtain enough statistics at DUNE for $E_{\nu}\lesssim 1$ GeV (around the 2nd oscillation maximum), good energy resolution (better than 10%), and near-detector measurements with a precision of order 1% or better.
@@ 2.1 Framework and assumptions
$\phi_{ij}\approx\frac{\Delta m^{2}_{ij,{\rm eff}}(E_{\nu})L}{4E_{\nu}}\,$, (2.11)
@@ 3.1 Experimental setups
DUNE is an upcoming long-baseline accelerator neutrino experiment designed to explore the nature of neutrinos by sending them from Fermilab to a far-detector situated deep underground at the Sanford Lab in South Dakota at a distance of 1300 km. To estimate the experiment’s sensitivity, we utilize the configurations outlined in the Technical Design Report (TDR) [25, 34], along with an alternative configuration featuring enhanced energy resolution as suggested in Refs. [35, 36] (see also [37, 38]). The TDR configuration includes a 40 kt Liquid Argon Time Projection Chamber (LArTPC) as the far detector (current planning considers a staged approach with up to 4 detector modules of 17 kt each), paired with a 120 GeV proton beam delivering 1.2 MW of beam power, which translates to $1.1\times 10^{21}$ protons on target (P.O.T.) per year. For further details on systematic errors and efficiencies, please refer to Ref. [34].
T2HK (Tokai to Hyper-Kamiokande) is an off-axis, accelerator-based future superbeam experiment with a 295 km baseline. To estimate the detector’s physics potential, we adhere to the experimental configurations outlined in [24]. This experiment will utilize the same 30 GeV proton beam from the J-PARC facility, previously used for T2K, to generate (anti)neutrino fluxes. The Water Cherenkov far detector is expected to have a fiducial volume of 187 kt, and the total exposure will be $1.3\,\rm{MW}\times 10\times 10^{7}$ seconds, equivalent to $2.7\times 10^{22}$ protons on target (P.O.T.). In this simplified scenario, we consider an uncorrelated 5% (3.5%) signal normalization error, a 10% background normalization error, and a 5% energy calibration error for both $\nu$ and $\bar{\nu}$ appearance (disappearance) channels.
@@ 3.2 Analysis details
[…]
@@ 4.1 Exposure and energy resolution
Let us now study the interplay of energy resolution and exposure. The estimates in section 3.3 suggest that the statistical uncertainty is dominated by DUNE under our default assumptions of 608 (840) kt MW yr for T2HK (DUNE), see eq. 3.5. Figure 4 shows the T violation sensitivity for $\delta_{\rm CP}^{\rm tr}=90^{\circ}$ as a function of the DUNE exposure in neutrino mode for different assumptions on the energy resolutions in both experiments. We see that an improved energy reconstruction can somewhat reduce the required exposure to reach a certain sensitivity. Note that our default assumption for DUNE according to eq. 4.1 corresponds already to an improved resolution according to [36]. The magenta curve in fig. 4 shows the sensitivity for a DUNE resolution reduced by a factor two compared to eq. 4.1, which corresponds to good accuracy to the value from the DUNE TDR [25, 34]. We see that the improved reconstruction is essential to obtain good sensitivity with reasonable exposures.
@@ 4.2 Zero-distance effect and prior on oscillation frequencies
Figure 5 shows a non-trivial interplay of the near-detector constrains (curves with different colors), the energy resolution (solid curves: default assumption, dashed curves: improved by a factor 0.5), and the oscillation frequency prior. A crucial role is played by the quantity $\delta_{0}$ defined in eq. 2.18, which multiplies the non-unitarity factor $\epsilon^{2}$ in the observable $X_{T}$, see eq. 2.17. The value of $\delta_{0}$ is determined by the oscillation frequencies and therefore depends on energy. Hence, effectively we have to average $\delta_{0}$ over the width of the considered energy bin and fold it with the resolution function. It turns out that the energy-averaged value of $\delta_{0}$ monotonically decreases from about 0.09-0.09 for our default energy resolution to 0.21-0.21 for a resolution improved by a factor 0.2. Hence, we expect that the impact of the zero-distance effect becomes more important for better energy resolution. This is the behaviour visible in fig. 5: while the solid curves are rather insensitive to the value of the near-detector constraint even up to $\sigma_{\epsilon}=5\%$, we see a rather strong dependence on $\sigma_{\epsilon}$ for the improved energy resolution (dashed curves) and the sensitivity degrades significantly for uncertainties $\sigma_{\epsilon}\gtrsim 1\%$. This behaviour is consistent with the estimate in eq. 3.5. Note that for $\sigma_{\epsilon}=5\%$ the better resolution leads to a worse sensitivity, which is a manifestation of the larger value of $|\delta_{0}|$ due to the non-trivial energy averaging, see also eq. 3.4.
@@ 5 Summary and discussion
Let us summarize in which sense our approach is model-independent: our analysis is very general in terms of the effective mixing parameters relating the flavour states and the eigenstates of the propagation Hamiltonian. We treat the $c_{i}$ coefficients (see eqs. 2.3 and 2.2) completely unconstrained in each energy bin and separate for neutrinos and antineutrinos. The strongest assumption is that only two oscillation frequencies are relevant, and that they are numerically close (within $\lesssim 2\%)$ to the standard oscillation frequencies in matter assuming the Standard Model matter effect. Note that this assumptions does not only involve the (rather precisely measured) vacuum mass-squared differences but also the impact of non-standard interactions potentially modifying the matter effect. The oscillation frequencies determine the conditions eqs. 2.19 and 2.20 on the neutrino energy and the two baselines which have to be fulfilled in order for the observable $X_{T}$ being sensitive to T violation. Sizeable uncertainties in the frequencies lead to a complicated interplay of energy resolution and zero-distance effects, affecting the sensitivity to T violation in a non-trivial way, see discussion in section 4.2.
@@ Appendix A Time reversal in the QFT formalism for neutrino oscillations
$P_{\alpha\to\beta}(L)$ $\propto\overline{|\mathcal{A}_{\alpha\to\beta}^{\rm red}(T,L)|^{2}}\propto\int dT|\mathcal{A}_{\alpha\to\beta}^{\rm red}(T,L)|^{2}$ (A.6)
$\propto\sum_{jk}U_{\alpha j}^{*}U_{\beta j}U_{\alpha k}U_{\beta k}^{*}\int dp^{0}\exp\left[i\frac{\Delta m^{2}_{kj}L}{2p^{0}}-f_{j}(p^{0})-f_{k}(p^{0})\right]$ (A.7)
@@ Appendix B Constraints on the zero-distance effect
    where NN ($D$) is the numerator (denominator) of eq. B.3. In the limit of small $\sigma_{\epsilon,\rm ext}$ we obtain